OpenStudy (anonymous):

How can you get coefficients that are equal for x? 12x+6y=13 4x+6y=12

OpenStudy (anonymous):

Is the question to solve for x?

OpenStudy (texaschic101):

I think that is what it is basically saying...solving for x

OpenStudy (anonymous):

no just asking How can you get coefficients that are equal for x? 12x+6y=13 4x+6y=12 A. Multiply both sides of the bottom equation by 2 B. Divide both sides of the bottom equation by 3 C. Divide both sides of the bottom equation by 2 D. Multiply both sides of the bottom equation by 3

OpenStudy (texaschic101):

I would multiply either equation by -1, then use elimination

OpenStudy (texaschic101):

None of your answer choices make much sense

OpenStudy (texaschic101):

Is the one that starts with 4x the bottom equation ?

OpenStudy (anonymous):

yea

OpenStudy (texaschic101):

@gypsy1274 ...want to take a look at this....maybe you can make some sense of it

OpenStudy (anonymous):

messgae her ?

OpenStudy (texaschic101):

@AccessDenied .....a little help here please

OpenStudy (texaschic101):

@Loser66 ...does this make sense to you ?

OpenStudy (loser66):

D

OpenStudy (texaschic101):

multiply by 3 12x + 18y = 36......I don't understand what that will do

OpenStudy (loser66):

hey, what is the question? " make it the same coefficient for x ". that's it. they don't ask you to solve . they don't ask you whether it makes sense or not, they just want to know by how you make 2 equations has the same coefficient of x. THAT' S IT.

OpenStudy (texaschic101):

oh....I misunderstood the question....I thought they were asking what x was....solving for x.....thank you so much Loser

OpenStudy (loser66):

yw

OpenStudy (texaschic101):

there you go ineedhelp....Loser got it for you

OpenStudy (anonymous):

@ineedhelp15 Do you understand the answer?

OpenStudy (texaschic101):

I really feel stupid on this one gypsy....lol

OpenStudy (anonymous):

No need to feel stupid. You should see the circles I had one woman going in last night, or the comedy of errors on the question myself and about 5 others just finished a few minutes ago... :(

OpenStudy (texaschic101):

well...ineedhelp finally got the answer...lol

OpenStudy (anonymous):

That's good. I'm going to log off for a few...I think I need to restart my computer. The delete button doesn't work, but only in the web browser. It works fine in all other programs.

OpenStudy (anonymous):

thanks every for your help i was stuck on it for days lol ...

OpenStudy (anonymous):

I'm glad you took time to think about it and try before just asking. You will learn more that way.